OpenStudy (jotopia34):

is the integral of x/x^2+3 equal to 1/2ln[x^2+3]+C?

OpenStudy (slaaibak):

Using the substitution x^2 + 3 = u du = 2x dx xdx = 1/2 du so it reduces to 1/2 int {1/u} du = 1/2 ln|x^2 + 3| + C so yes, you're correct :)

OpenStudy (jotopia34):

YAYYYY, did that one myself...hoowah

OpenStudy (slaaibak):

well done, i'm proud :)